# The Geometry behind the Two-Dimensional Hard-Disk ‘Glass Transition’

### v1

@@ I.5 Structure of the Paper
The remainder of this paper is organized as follows. Section II provides the details necessary to reproduce my results and Section III the ways, in which I characterized the structure in the bidisperse hard-disk system. The major features of working with the isoperimetric quotient are put forward in Section IV, which is followed by an analysis of the dynamics in the system via the MSD in Section V. Together, these results lead me to propose a state diagram in Section VI, which is discussed in detail. Next, I correlate structural features to particle dynamics in Section VII, which is supplemented by a literature analysis (Appendix A). Then, I present my geometric model and understanding in Section VIII. The implications of the study are further contextualized in the discussion, where I also give a perspective on an extension to 3D, see Section IX and Appendix F. Lastly, I provide a short summary of the work in Section X.
@@ II System and Numerical Approach
I study structural and kinetic properties of systems of binary hard disks in 2D using the EDMD software developed by Smallenburg [71]. The algorithms were appropriately modified for use in 2D, see the ESI [59]. In this variant of EDMD, the system is thermalized and maintained at a reduced temperature of $T=1$. The choice of parameters is such that in the limit of $\phi\downarrow 0$ the diffusion coefficient of the particles is given by $D=1$. That is, the reduced time is expressed in the diffusion time.
The investigated systems consist of a 50:50 mixture of small and large discs, for which two particles (indexed $i$ and $j$) interact via the pair-interaction potential
$U(r)$ =$=\left\{\begin{array}[]{cc}0&2r>\sigma_{i}+\sigma_{j}\\ \infty&2r\leq\sigma_{i}+\sigma_{j}\end{array}\right.,$ (4)
@@
In all cases, I simulated $N=$ 2,048 particles in a square simulation box with edge length $L$ and periodic boundary conditions  [^4] I performed a small number of simulations with 10,000 particles, which showed little difference to the results obtained for a smaller number of particles.. The value of $L$ was varied such that I could simulate area fractions $\phi=(\pi/8)\sigma^{2}\left(1+R^{2}\right)N/L^{2}$ between 00 and $\phi_{m}$. The latter is the random-close packing (rcp) area fraction, $\phi_{m}$. I established $\phi_{m}$ for a given $R$ using the approach of Desmond and Weeks [75], the limitations of which are discussed in Appendix A. In brief, I computed 250 potential (random) closest packings using the same number of particles $N$ as employed for my production runs. I then selected the highest value of $\phi$ and thus obtained the rcp area fraction $\phi_{m}$ as a function of $R$, see Section VI for the result.
I typically considered 100+ non-equidistant points $\phi\in[0,\phi_{m}]$ for each value of $R$. For each $R$-$\phi$ combination, I then generated 150 independent realizations of the system by slowly growing the particles until the target value of $\phi$ had been obtained, see the ESI [59]. A dimensionless growth rate of $10^{-3}$ proved sufficient to obtain initial configurations for all $\phi$ even very close to $\phi_{m}$. It should be noted, however, that the time for finding an initial configuration increased substantially as $\phi\uparrow\phi_{m}$, which is expected for the highly constrained packing.
Next, I allowed each configuration to evolve for $10^{4}$ reduced time units without growth, before I started to sample. As we will see, there are slow dynamics in the system and $10^{4}$ time units may be too short of a time to fully equilibrate the system for all considered $\phi$-$R$ combinations. Here, I focused on removing any bias that could be present in the system at low to intermediate $\phi$ by creating the configurations via growth. The 150 configurations per $R$-$\phi$ point were used to obtain error bars and relatively accurate statistics.
Lastly, for the purposes of completeness, I should note that in addition to EDMD [71], I have attempted to create and equilibrate systems using plain Metropolis, rejection-free [76, 77], and swap variants of the Monte Carlo algorithm [78, 79, 21]. However, none of these methods proved to be more efficient than the EDMD algorithm used here. In particular, the use of swap Monte Carlo proved challenging, due to the bidisperse hard-particle nature of the sample at small values of $R$. That is, most moves were rejected beyond a certain value of $\phi$. Though, I should note that by no means this implies that these algorithms are not generally useful to study hard-particle systems with arrested kinetics, for example, as recently evidenced in Ref. [80].
@@ III Characterization
The purpose of this paper is to probe the connection between arrested dynamics and isoperimetric properties of the local particle neighborhoods. I therefore compute the radical Voronoi tessellation (also known as power diagram or Dirichlet cell complex) using the voro++ library [81, 82]. The radical Voronoi tessellation respects the radius difference between the two sizes of particles in the system. Figure 1a shows a result of this analysis for a small part of a simulation volume in for $R^{-1}=1.4$ at an area fraction of $\phi=0.777$; close to the experimental value for the glass transition [1].
Figure 1: Procedure for computing the isoperimetric quotient using a Voronoi tessellation. (a) Part of a snapshot of a system showing the instantaneous organization of bidisperse disks with size ratio $R^{-1}=1.4$ at an area fraction of $\phi=0.777$. The smaller disks are indicated in light red, while the larger disks are in the darker shade of red. The centers of the disks are indicated using black dots and the blue lines surrounding the disks represent the radical Voronoi diagram for the system. Note that none of the disks overlap though this is hard to resolve by eye. (b) An example of a Voronoi cell for a large particle which has 6 sides, which is the $i$-th cell in this sample. (c) The area $A_{i}$ of this cell (green) together with the perimeter length $P_{i}=\sum_{n}l_{i}^{n}$ (purple arrows) are used to compute the isoperimetric quotient $q_{i}=4\pi A_{i}/P_{i}^{2}$.
Using the Voronoi cells obtained in this manner, I compute the isoperimetric quotient
@@
for each particle, where $A$ is the area of the associated Voronoi neighborhood and $P$ is its perimeter. Note that $q\in(0,1]$, where $q=0$ corresponds to an extremely needle-like Voronoi neighborhood, while $q=1$ represents a perfect circle. Recall that for each $R$-$\phi$ combination, there are 150×150$\times$150 ×2,048 $q$ values. This allows me to compute using non-linear binning the probability density function for $q$, here denoted $P_{\phi}(q)$, to a high degree of accuracy, as I will present in Section IV.
The distribution $P_{\phi}(q)$ is insightful, but to help understand the properties of the underlying system, it is beneficial to convert $q$ to an effective number of vertices $n_{q}$ [83]. The isoperimetric quotient of regular polygons with $n$ edges can be readily computed and reads
@@
where $A_{\mathrm{r}}(n)$ and $P_{\mathrm{r}}(n)$ are the area and the perimeter of a regular polygon having $n$ edges and a circumscribing circle with radius $r=1$. This function maps the number of the edges of regular polygons to their isoperimetric quotient and vice versa. The right-hand side of Eq. (6) can now straightforwardly be applied to non-integer values of $n$. I introduce the concept of a generalized number of edges to underscore that non-integer values are permissible, such as $n=5.2$. In addition, I make use of the subscript ‘$q$’ to indicate the origin of this number: $n_{q}$.
Numerically inverting Eq. (6), I am able to obtain an $n_{q}$ for any value of $q$. Note that there is only a single polyhedron that maps to integer values of $n_{q}$, namely the regular one. However, there are many, subtly different polyhedra that give, for example, the value $n_{q}=5.2$. The point of introducing the generalized number of edges is to see all of these as roughly the same and close to pentagonal, as already though less clearly expressed by $q$. The definition of $n_{q}$ also allows me to identify ranges in $q$ that intuitively correspond to certain polygons [83]. That is, I assign a specific value of $n_{q}$ to belong to triangular, square, pentagonal, hexagonal, heptagonal, etc. neighborhoods by taking $\lfloor n_{q}\rceil$. This choice factors the $n_{q}$ (equivalently $q$) range into distinct subsets $n$-fold pseudo polygons. For example, all neighborhoods having $\lfloor n_{q}\rceil=5$ (i.e., $4.5<n_{q}\leq 5.5$) can be referred to as pseudo-pentagons. Integrating the PDF belonging to the pseudo-pentagons provides insight into the fraction of pentagonal neighborhoods in the system, also see Appendix B. This includes neighborhoods with more than five edges, for which some edges are very small and therefore the neighborhood has the appearance of a pentagon, Fig. 1 shows several examples of this.
I also computed several ensemble-averaged and time-dependent physical quantities of interest to further support my arguments. These include the pressure $p$, see Appendix C, and two variants of cluster percolation, see Appendix D. Lastly, I determined the kinetics in the bidisperse hard-disk samples by computing the MSD. I obtained particle displacements for $10^{4}$ unit times starting from the ‘equilibrated’ samples. These were averaged in time, which led to 150 MSD curves that were also averaged over all particles in the system. Subsequently, I determined the mean of the independent MSD data — the initial configuration varies — as well as the associated standard deviation. Section V provides additional details on the processing of this data, by which I obtained the area fraction $\phi_{a}$ for which arrested dynamics sets in.
@@ IV Neighborhood Structure
Figure 2: Structural features of the 2D system characterized by the isoperimetric quotient. (a-c) The probability density function (PDF) $P_{\phi}(q)$ of the isoperimetric quotient $q$ for several values of the area fraction $\phi$ in lin-log representation. The inverse size ratio $R^{-1}$ between the large and small spheres in the system is provided in by the label. From red to purple, the area fractions are $\phi=0.350$ (red), $0.500$ (orange), $0.610$ (green), $0.720$ (cyan), $0.780$ (blue), $0.825$ (purple), respectively. The vertical black lines provide the $q$ values for a regular $n$-gon, with $n=3,\dots,7$ from left to right. The error bars indicate the standard deviation on the binned value for the interval over which $P_{\phi}(q)$ is sampled, which adheres to counting statistics and is thus always the square root of the number of samples in a bin. (d) The generalized number of edges $n_{\bar{q}}$ belonging to the mean $q$ value — $\bar{q}$: the first moment of $P_{\phi}(q)$ — as a function of $\phi$ for all size ratios $R$ considered. From red to purple the value of the size ratio is given by $R^{-1}=1.00$, $1.05$, $1.10$, $1.15$, $1.17$, $1.20$, $1.22$, $1.23$, $1.24$, $1.25$, $1.30$, $1.35$, $1.40$, $1.45$, $1.50$, $1.55$, $1.60$, $1.65$, and $1.70$, respectively. The inset shows a zoom-in on the range where the systems form crystals, i.e., above the dashed horizontal line provides the value $n_{\bar{q}}=5.5$. I do not provide error bars here to help improve the representation. The error is about twice the line width, as can be appreciated from the noise that is present on the red curve in the inset.
Now let us turn to the outcome of my structural characterization. Figure 2a-c shows a sampling of the isoperimetric-quotient PDFs, $P_{\phi}(q)$, for three values of $R$ and 6 values of $\phi$, as labelled. Note that irrespective of $R$, the low-$\phi$ diagram is similar. The distribution tends to an established curve as $\phi\downarrow 0$ that was computed for Voronoi diagrams deriving from uniformly distributed points [84, 85]. This makes sense, as for low $\phi$, properties such as differences in particle size should wash out of the distributions, as also seen in the literature [86].
@@
Comparing the high-$\phi$ PDFs for the three $R$ values, I observe that the monodisperse ($R=1$) system convincingly crystallizes. The PDF has a single peak which is close to $q_{\mathrm{r}}(6)$, as expected. For lower values of $R$, there remains a sizable fraction of pentagons and there can also be a significant number of heptagons, when $R$ is sufficiently small. I refer to Appendix B for additional details on the fraction of pseudo-polygons. For systems with $R^{-1}\gtrsim 1.4$, I consistently find that the two peaks in the ranges $(q_{\mathrm{r}}(4),q_{\mathrm{r}}(5)]$ and $(q_{\mathrm{r}}(5),q_{\mathrm{r}}(6)]$ are comparable in height. However, there is a slight abundance of pseudo-pentagons, see Appendix B. This means that the disk configurations are not crystalline.
The features of the $P_{\phi}(q)$ will be further analyzed in Section VII. However, before turning to dynamics in Section V, I will first consider the mean value of $q$ that can be obtained from $P_{\phi}(q)$ by taking the first mode of the distribution. Let this be denoted by $\bar{q}$ and its associated generalized number of edges by $n_{\bar{q}}$. The latter is shown as a function of $\phi$ in Fig. 2d for all values of $R$ considered. The curves are generally smooth within the error. There is no apparent gap or discontinuity for those systems that are known to undergo a phase transition. This is because I worked in the constant area, particle, and temperature ensemble. There appears to be a change in slope close to where the system is expected to crystallize (for $R\lesssim 1$), but I did not investigate this further. The generalized number of edges $n_{\bar{q}}$ is an increasing function of $\phi$ for $\phi\lesssim 0.76$. For greater area fractions, I find a single, well-defined maximum, beyond which $n_{\bar{q}}$ decreases rapidly. I will use this feature, as well as the intersection points with $n_{\bar{q}}=5.5$, see inset to Fig. 2d, in preparing a $R$-$\phi$ state diagram in Section VI.

### v2

@@ I.6 Structure of the Paper
The remainder of this paper is organized as follows. Section II provides the details necessary to reproduce my results and Section III the methods by which I characterized the structure in the bidisperse hard-disk system. The major features of working with the isoperimetric quotient $q$, which is a variant of the aforementioned shape index $p$ ($q=4\pi/p^{2}$), are put forward in Section IV. This is followed by an analysis of the dynamics in the system via the MSD in Section V. Together, these results lead me to propose a state diagram in Section VI, which is discussed in detail. Next, I correlate structural features to particle dynamics in Section VII, which is supplemented by a literature analysis (Appendix A). Then, I present my geometric model and understanding in Section VIII. The implications of the study are further contextualized in the discussion, where I also give a perspective on an extension to 3D, see Section IX and Appendix G. Lastly, I provide a short summary of the work in Section X.
@@ II System and Numerical Approach
I study structural and kinetic properties of systems of binary hard disks in 2D using the EDMD software developed by Smallenburg [100]. The algorithms were appropriately modified for use in 2D, see the ESI [64]. This variant of EDMD works in the canonical ($NAT$) ensemble  [^5] Variants of EDMD that solve for Brownian dynamics are possible [152]. Here, the thermalization is introduced by redrawing the velocity of a random selection of particles from the Maxwell-Boltzmann distribution.. The choice of parameters is such that in the limit of $\phi\downarrow 0$ the (extrapolated) diffusion coefficient of the particles is (typically) given by $D=1$  [^6] Note that the collision rate for small $\phi$ is itself small, given the low density of particles. As such, the algorithm does not reproduce diffusion correctly in the limit $\phi\downarrow 0$, which can result in a strong increase of the diffusion coefficient at low $\phi$. This effect is present is in some of the data, but not shown here.. That is, the reduced time is expressed in terms of the bare diffusion time and I verified that this was the case.
The investigated systems consist of a 1:1 mixture of small and large discs, for which two particles (indexed $i$ and $j$) interact via the pair-interaction potential
$U(r)$ $=\left\{\begin{array}[]{cc}0&2r>\sigma_{i}+\sigma_{j}\\ \infty&2r\leq\sigma_{i}+\sigma_{j}\end{array}\right$., (4)
@@
In all cases, I simulated $N=$ 2,048 particles in a square simulation box with edge length $L$ and periodic boundary conditions  [^7] I performed a small number of simulations with 10,000 particles, which showed little difference to the results obtained for a smaller number of particles.. The value of $L$ was varied such that I could simulate area fractions $\phi=(\pi/8)\sigma^{2}\left(1+R^{2}\right)N/L^{2}$ between 0 and $\phi_{m}$. Here, $\phi_{m}$ is the RCP area fraction, which I established for a given $R$ using the approach of Desmond and Weeks [109], the limitations of which are discussed in Appendix A. In brief, I computed 250 potential (random) closest packings using the same number of particles $N$ as employed for my production runs. From these I selected the highest value of $\phi$ and thereby obtained an estimate for the RCP area fraction $\phi_{m}$ as a function of $R$, see Section VI for the result.
I typically considered 100+ non-equidistant points $\phi~\in~[0,\phi_{m}]$ for each value of $R$. For each $R$-$\phi$ combination, I generated 350+ independent realizations of the system by slowly growing the particles until the target value of $\phi$ had been obtained, see the ESI [64]. A dimensionless growth rate of $10^{-3}$ proved sufficient to obtain initial configurations for all $\phi$ even very close to $\phi_{m}$. It should be noted, however, that the time for finding an initial configuration increased substantially as $\phi\uparrow\phi_{m}$, which is expected for the highly constrained packing.
Next, I allowed each configuration to evolve for $10^{4}$ reduced time units without growth, before I started to sample. It will become clear that these $10^{4}$ time units proved sufficient to equilibrate the system for most considered $\phi$-$R$ combinations, despite the slow dynamics. This time only proved insufficient very close to RCP at low $R$ values, see ESI [64], which are not relevant for the conclusions drawn in this study  [^8] For the purposes of completeness, I should note that in addition to EDMD [100], I have attempted to create and equilibrate systems using plain Metropolis, rejection-free [153, 154], and swap variants of the Monte Carlo algorithm [155, 156, 25]. However, none of these methods proved to be more efficient than the EDMD algorithm used here. In particular, the use of swap Monte Carlo proved challenging, due to the bidisperse hard-particle nature of the sample at small values of $R$. That is, most moves were rejected beyond a certain value of $\phi$. Though, I should note that by no means this implies that these algorithms are not generally useful to study hard-particle systems with arrested kinetics, for example, as recently evidenced in Ref. [157]..
@@ III Characterization
I probed structure by computing the radical Voronoi tessellation (also known as power diagram or Dirichlet cell complex) using the voro++ library [111, 112]. The radical Voronoi tessellation respects the radius difference between the two sizes of particles in the system. Figure 3a shows a result of this analysis for a small part of a simulation volume in for $R^{-1}=1.4$ at an area fraction of $\phi=0.777$; close to the experimentally relevant value of the area fraction [82, 1, 3].
Figure 3: Procedure for computing the isoperimetric quotient using a Voronoi tessellation. (a) Part of a snapshot of a system showing the instantaneous organization of bidisperse disks with size ratio $R^{-1}=1.4$ at an area fraction of $\phi=0.777$. The smaller disks are indicated in light red, while the larger disks are in the darker shade of red. The centers of the disks are indicated using black dots and the blue lines surrounding the disks represent the radical Voronoi diagram for the system. Note that none of the disks overlap though this is hard to resolve by eye. (b) An example of a Voronoi cell for a large particle which has $n=6$ sides, which is the $i$-th cell in this sample. (c) The area $A_{i}$ of this cell (green) together with the perimeter length $P_{i}=\sum_{i=1}^{n}l_{i}$ (purple arrows) are used to compute the isoperimetric quotient $q_{i}=4\pi A_{i}/P_{i}^{2}$.
Using the Voronoi cells obtained in this manner, I computed the isoperimetric quotient
@@
for each particle, where $A$ is the area of the associated Voronoi neighborhood and $P$ is its perimeter. Note that $q\in(0,1]$, where $q=0$ corresponds to an extremely needle-like Voronoi neighborhood, while $q=1$ represents a perfect circle. Recall that for each $R$-$\phi$ combination, there are (typically) $350\times$2,048 $q$ values. This allows me to compute the probability density function (PDF), here denoted $P_{\phi}(q)$, using non-linear binning for $q$ to a high degree of accuracy, as I will present in Section IV.
Importantly, the choice for $q$ rather than $p$ is meaningful, because for specific congruent tilings of the plane, the area fraction of monodisperse disks inscribed in this tiling equals the value of $q$. That is, $\phi=q$ for these special configurations. The proof is as follows: Take a tangential polygon — this has an inscribed circe (diameter $\sigma$) touching all sides — with $n$ sides of potentially different length. The area of this disk is then $\pi\sigma^{2}/4$ and the polygon’s area can be computed by dividing it into $n$ triangles by drawing line segments from the incenter to each vertex. For each side $l_{i}$ the triangle’s surface area is $\sigma l_{i}/4$. Therefore, the total surface area is $A=(\sigma/4)\sum_{i=1}^{n}l_{i}=\sigma P/4$. Hence, the area fraction is $\phi=(\pi\sigma^{2}/4)/(\sigma P/4)=\pi\sigma/P$. But, from the area, it follows $\sigma=4A/P$ which leads to $\phi=4\pi A/P^{2}=q$. The interpretation useful to this work is that if a constant value of $\phi$ is found, there is an automatic connection to a congruent tiling of the plane. For this tiling all particles must touch, such that the radical Voronoi diagram consists of tangential polygons only.
To improve understanding — $n=5$ edges is more transparent than $q\approx 0.865$ — it proves beneficial to convert $q$ to a generalized number of edges $n_{q}$ [113], here the subscript “qq” is used to indicate the origin of this number. The isoperimetric quotient of regular polygons with $n$ edges is unique and reads
@@
where $A_{\mathrm{r}}(n)$ and $P_{\mathrm{r}}(n)$ are the area and the perimeter of a regular polygon having $n$ edges and a circumscribing circle with radius $r=1$. Numerically inverting Eq. (6), I am able to obtain an $n_{q}$ for any value of $q$, where now non-integer values of the edge number are permissible, such as $n_{q}=5.2$  [^9] Note that there is only a single polyhedron that maps to integer values of $n_{q}$, namely the regular one. However, there are many, subtly different polyhedra that give, for example, the value $n_{q}=5.2$. The point of introducing the generalized number of edges is to see all of these as roughly the same and close to pentagonal.. Additionally, the definition of $n_{q}$ also allows me to identify ranges in $q$ belonging to triangular, square, etc. neighborhoods by taking $\lfloor n_{q}\rceil$. For example, all neighborhoods having $\lfloor n_{q}\rceil=5$ (i.e., $4.5<n_{q}\leq 5.5$) can be referred to as pseudo-pentagons. Integrating the PDF belonging to the pseudo-pentagons provides insight into the fraction of pentagonal neighborhoods in the system, also see Appendix B.
To complement this Voronoi-based analysis, I also computed the traditional six-fold Steinhardt order parameter $\psi_{6}$, which has the form
$\psi_{6}$ $=\left\langle\frac{1}{N}\textstyle\sum_{i=1}^{N}\left|\textstyle\sum_{j=1}^{N_{n}(i)}\frac{l_{i}^{j}}{P_{i}}\exp\left(6\iota\theta_{ij}\right)\right|\right\rangle.$ (7)
[…]
I also computed several ensemble-averaged and time-dependent physical quantities of interest to further support my arguments. These include the pressure $p$, see Appendix C, and two variants of cluster percolation, see Appendix D. Lastly, I determined the kinetics in the bidisperse hard-disk samples by computing the MSD. I obtained particle displacements for $10^{4}$ unit times starting from the equilibrated samples. This led to 150 MSD independent curves that were averaged over all particles in the system. Subsequently, I determined the mean of the independent MSD data — the initial configuration varies — as well as the associated standard deviation. Section V provides additional details on the processing of this data, by which I obtained the area fraction $\phi_{a}$ for which the diffusion vanishes.
@@ IV Neighborhood Structure
Figure 4: Structural features of the 2D system characterized by the isoperimetric quotient. (a-c) The probability density function (PDF) $P_{\phi}(q)$ of the isoperimetric quotient $q$ for several values of the area fraction $\phi$ in lin-log representation. The inverse size ratio $R^{-1}$ between the large and small disks in the system is provided in by the label. From red to purple, the area fractions are $\phi=0.351$, $0.341$, $0.339$ (a-c; red), $0.503$, $0.500$, $0.497$ (a-c; orange), $0.616$, $0.608$, $0.606$ (a-c; green), $0.720$ (cyan), $0.780$ (blue), $0.825$ (purple), respectively. The vertical black lines provide the $q$ values for a regular $n$-gon, with $n=3,\dots,7$ from left to right; also see visual guide on top. The error bars indicate the standard deviation on the binned value for the interval over which $P_{\phi}(q)$ is sampled, which adheres to counting statistics and is thus always the square root of the number of samples in a bin. (d) The generalized number of edges $n_{\bar{q}}$ belonging to the mean $q$ value — $\bar{q}$: the first moment of $P_{\phi}(q)$ — as a function of $\phi$ for all size ratios $R$ considered. From red to purple the value of the size ratio is given by $R^{-1}=1.00$, $1.05$, $1.10$, $1.15$, $1.17$, $1.20$, $1.22$, $1.23$, $1.24$, $1.25$, $1.30$, $1.35$, $1.40$, $1.45$, $1.50$, $1.55$, $1.60$, $1.65$, and $1.70$, respectively. The inset shows a zoom-in on the range where the systems form crystals, i.e., above the dashed horizontal line provides the value $n_{\bar{q}}=5.5$. I do not provide error bars here to help improve the representation. The error is about twice the line width, as can be appreciated from the noise that is present on the red curve in the inset.
Figure 4a-c shows a sampling of the isoperimetric-quotient PDFs, $P_{\phi}(q)$, for three values of $R$ and six values of $\phi$, as labelled. Note that irrespective of $R$, the low-$\phi$ diagram is similar. The distribution tends to an established curve as $\phi\downarrow 0$ that was computed for Voronoi diagrams deriving from uniformly distributed points [115, 116]. This makes sense, as for low $\phi$, properties such as differences in particle size should wash out of the distributions, as also seen in the literature [117].
@@
Comparing the high-$\phi$ PDFs for the three $R$ values, I observe that the monodisperse ($R=1$) system convincingly crystallizes. The PDF has a single peak which is close to $q_{\mathrm{r}}(6)$, as expected. For lower values of $R$, there remains a sizable fraction of pentagons and there can also be a significant number of heptagons, when $R$ is sufficiently small. I refer to Appendix B for additional details on the fraction of pseudo-polygons. For systems with $R^{-1}\gtrsim 1.4$, I consistently find that the two peaks in the ranges $(q_{\mathrm{r}}(4),q_{\mathrm{r}}(5)]$ and $(q_{\mathrm{r}}(5),q_{\mathrm{r}}(6)]$ are comparable in height. However, there is a slight abundance of pseudo-pentagons, also see Appendix B. This means that the disk configurations are not crystalline.
The features of the $P_{\phi}(q)$ will be further analyzed in Section VII, where we will connect them to $\psi_{6}$. However, before turning to dynamics in Section V, I will first consider the mean value of $q$ that can be obtained from $P_{\phi}(q)$ by taking the first mode of the distribution. Let this be denoted by $\bar{q}$ and its associated generalized number of edges by $n_{\bar{q}}$. The latter is shown as a function of $\phi$ in Fig. 4d for all values of $R$ considered. The curves are generally smooth within the error. There is no apparent gap or discontinuity for those systems that are known to undergo a phase transition. This is because I worked in the canonical ensemble and used a limited number of particles. There appears to be a change in slope close to where the system is expected to crystallize (for $R\lesssim 1$), but I did not investigate this further. The generalized number of edges $n_{\bar{q}}$ is an increasing function of $\phi$ for $\phi\lesssim 0.76$. For greater area fractions, I find a single, well-defined maximum, beyond which $n_{\bar{q}}$ decreases rapidly. I will use this feature, as well as the intersection points with $n_{\bar{q}}=5.5$, see inset to Fig. 4d, in preparing a $R$-$\phi$ state diagram in Section VI.